# Quantifying Resolution Limits in Pedestal Profile Measurements with Gaussian Process Regression

### v1

@@ 5 Application to DIII-D Pedestal Data
In this section, we focus on applying the developed techniques to fit pedestal profiles from the DIII-D tokamak [40], discharge #174864. This discharge, which comes from a series of discharges studying the physics of microtearing modes (MTMs) in DIII-D [41], has a particularly wide pedestal as well as sufficient time resolution from the measurements to provide some information about the dynamics of the pedestal between edge localized modes (ELM). In section 5.1 we consider profiles depending on the radial coordinate only, while in sections 5.2 and 5.3 we see how the tools can be used to verify the credibility of features in time-dependent profiles.
@@ 5.1 ELM-averaged Fits
Figure 4: GPR inferences for the electron temperature $T_{e}$ as a function of normalized poloidal flux $\psi$ for DIII-D shot #174864. (Top and middle rows) Data (blue scatter) are shown with corresponding inferred profiles (orange with ±$\pm 34\%$ confidence interval shaded). The inverse filter cutoff $1/\xi_{*}$ is also shown. Comparing to the width of the pedestal, the right-most profile shows signs of over-regularization. (Bottom row) $N_{eff}$ for the derivative profiles is shown as a function of $\psi$, with $N_{eff}\lesssim 2$ indicating overfitting.
@@
We begin by considering the problem of inferring the profile of the electron temperature $T_{e}$ as a function of the normalized poloidal flux $\psi$ from measurements from the Thomson scattering system [42]. Data are aggregated from $N_{t}=44$ time slices spanning the ELM cycle, covering phases from 5% to 90% of the ELM cycle. The data, as well as GPR inferences for three different values of hyperparameters, are shown in figure 4.
@@
The leftmost column displays an inference based on hyperparameters $\sigma^{2},\ell$ set to their MLE values while keeping $\nu=2.5$ fixed. Some ‘wiggles’ appear in the inferred derivative profile, raising questions about the credibility of the inference. To quantify this, we consider $N_{eff}(\psi)=N_{\mathbf{Y}}[T_{e}^{\prime}(\psi)]$, which represents the effective number of measurements contributing to the inference of the profile derivative $T_{e}^{\prime}$ at a given radial location $\psi$. The bottom row of figure 4 plots this quantity normalized by the number of time slices $N_{t}$. The resulting normalized metric $N_{eff}(\psi)/N_{t}$ provides the effective number of measurements per time slice used to infer the derivative at $\psi$. $N_{eff}/N_{t}\lesssim 2$ indicates that the derivative is primarily inferred using ‘sub-channel’ resolution arising from the radial scatter of the measurements over different time slices. This radial scatter introduces potential covariances in the data that may not be fully captured by the assumed model for the measurement error $\varepsilon_{i}$, suggesting the MLE inference could be over-fitting fluctuations in the data in these regions.
@@
Figure 5: Plots showing how varying the length scale $\ell$ hyperparameter changes the inferred values for the peak gradient $\operatorname{max}|dT_{e}/d\psi|$ and full width at half maximum (FWHM) of the temperature pedestal. In addition, the bottom plot shows the inverse cutoff frequency $1/\xi_{*}$ compared to the FWHM. As $\ell$ increases so does $1/\xi_{*}$, leading to more strongly regularized inferences with decreased peak gradients and broader pedestal widths.
@@
The center and rightmost columns show inferences with increased $\ell$ relative to the MLE case, with $\nu,\sigma^{2}$ kept fixed. These hyperparameter values could plausibly arise from a hyperprior which biases the hyperparmeter estimator towards larger length scale parameters. Alternatively, they could arise as realizations of an ensemble of inferences in a Bayesian marginalization approach, as discussed in section 2.2. The metric $N_{eff}(\psi)/N_{t}\gtrsim 4$ in the steep gradient region, indicating that the inferred derivative values are indeed using information over multiple measurement channels, reducing concerns about overfitting sub-channel fluctuations in the data associated with the radial scatter. However, there are significant quantitative differences in the inferred profiles for these two cases – the peak gradients are 8% smaller, while the full width at half maximum (FWHM) $w_{ped}$ of the gradient profile is 13% larger for the rightmost case compared to the center case. This is suggestive of a certain amount of over-regularization causing the sharp gradient region to be smoothed out. Given that the physics of turbulent transport tends to suggest that profiles tend to stay clamped near marginal stability boundaries, such over-regularization could result in large differences in predicted mode stability or transport.
@@
In order to systematically quantify the level of over-regularization, we turn to the low-pass interpretation derived for GPR. We consider an effective $\overline{\Delta\psi}=0.2/N_{[0.8,1.0]}$, where $N_{[0.8,1.0]}$ is the total number of measurements of the interval $0.8\leq\psi\leq 1.0$, and $\overline{\sigma_{\varepsilon}^{2}}$ given by the average squared error of the data. We can then compute an effective cutoff frequency $\xi_{*}$ by using (7) with the average signal-to-noise rate $\bar{S}=\sigma^{2}\ell/\overline{\sigma_{\varepsilon}^{2}}\overline{\Delta\psi}$ In addition to plotting the inverse cutoff frequency $1/\xi_{*}$ in figure 4, we also show how the peak gradients, pedestal FWHM, and inverse cutoff frequency systematically change as we vary $\ell$ in figure 5. Note that when $1/\xi_{*}$ crosses the FWHM, the peak gradient and pedestal width become more sensitive to $\ell$, indicating the onset of over-regularization.
Taken together, the $N_{eff}$ metric and the low-pass filter interpretation lend a high degree of confidence to the credibility of the inference in the center column of figure 4. $N_{eff}/N_{t}\gtrsim 3$ provides confidence that the pedestal profile is not over-fit, while the low-pass cutoff frequency $\xi_{*}w_{ped}\lesssim 1$ provides confidence that the pedestal profile is not over-regularized.
@@ 5.2 Inter-ELM Variation: Setup
To accomplish this, we assume that the value of a measurement $Y_{i}$ depends on both the spatial and temporal location of its measurement, $x_{i}=(\psi_{i},t_{i})$. Thus, we are trying to infer a two-dimensional function $p(\psi,t)$ from noisy measurements $Y_{i}=p(\psi_{i},t_{i})+\varepsilon_{i}$. We take a GP prior on $p$ with zero mean, and a covariance which is the direct product of two kernels,
@@
We now use the tools developed in the previous sections to get a point estimate for a set of hyperparameters which provide a credible fit of the inter-ELM evolution. Since we are fitting a pedestal profile, we use $\nu_{\psi}=2.5$. Also, to capture sharp transitions in temporal behavior, we use $\nu_{t}=2.5$. We get an estimate for $\sigma^{2}$ by taking the mean square of the data $Y_{i}^{2}$. This leaves the two length scale parameters $\ell_{\psi},\ell_{t}$ to be determined.
@@
In figure 6, we use equation (8) to plot the transfer function $H(\xi,f)$ associated with the product kernel (17) over spatial $\xi$ and temporal $f$ frequencies for a representative choice of $\ell_{\psi}$ and $\ell_{t}$. Notice that the shape of the pass band enclosed by the $H(\xi,f)=1/2$ contour indicates that high temporal frequencies experience stronger attenuation in spatial frequencies, and vice versa.
@@
Next, we average the values of $N_{eff}$ over the spatial range $[0.8,1.0]$ and temporal range $[t_{min},t_{max}]$ of the data to compute average values $\bar{N}_{\mathbf{T}}$ and $\bar{N}_{\psi}$. These are used to define the effective spatial and temporal signal-to-noise rates
@@
By keeping the hyperparameters $\sigma^{2},\nu_{\psi},\nu_{t}$ fixed, we can plot $\bar{N}_{\mathbf{T}},\bar{N}_{\psi},\xi_{*}$ and $f_{*}$ as functions of $\ell_{\psi},\ell_{t}$. This is done in figure 7. The ELM period was about 100100~{}100100 ms in the considered time range, so a temporal cutoff frequency $f_{*}=30$ cycles per second was chosen, so phenomena with a timescale of $1/f_{*}\approx 33$ ms could be resolved. The spatial frequency $\xi_{*}$ was chosen so the product with the pedestal full width at half maximum $w_{ped}$, computed by fitting the profile at the given hyperparameter values, would satisfy $\xi_{*}w_{ped}=1$.
@@ 5.3 Inter-ELM Variation: Results
Figure 8: Time-dependent GPR inferences performed using the conditional averaging technique for the electron pressure profile $p_{e}=n_{e}T_{e}$ as a function of $\psi$. (Top and middle rows) Data (scatter) are shown with inferred profile and derivatives $dp_{e}/d\psi$ (curves with ±$\pm 34\%$ confidence intervals shaded), colored by the time since ELM. The inverse filter cutoff $1/\xi_{*}$ is shown along with low-order rational $q$ surfaces on the derivative profile. (Bottom row) $N_{\psi}$, averaged over $t$, is plotted as a function of $\psi$. $N_{\psi}\lesssim 2$ indicates possible overfitting due to reliance on two or less measurement channels per timeslice for the inference of the derivative.
@@
Recent work has shown how the alignment between the peak of the profile of electron diamagnetic frequency $\omega_{*}$ with low-order rational surfaces in the pedestal can be used to predict features in experimentally observed magnetic fluctuation spectrograms [43, 44, 45, 46, 47]. Computing profile gradients is crucial to such an analysis, as $\omega_{*}$ is related to pressure gradients via
@@
Furthermore, following the analysis [41], we are able to compare these pressure profiles against the locations of low-order rational $q$ surfaces. Namely, kinetic equilibria are produced every $20$ ms using the CAKE code [48], then sorted by the time since ELM. The locations of low-order rational $q$ surfaces are computed by averaging over their locations determined from the individual equilibria, and errorbars are computed using the standard error of the mean.
@@
The second interesting feature is the inward motion of the peak of the pedestal pressure gradient as the pedestal profile evolves during the inter-ELM cycle. In [41], this behavior was linked to up-chirping of magnetic fluctuations observed in the pedestal region through the predicted dependence of the MTM frequency on the peak value of $\omega_{*}$. We repeat the time-dependent MTM analysis using the profiles inferred from GPR. We plot the evolution of the peak of $T_{e}\frac{\partial(\log p_{e})}{\partial\psi}\propto\omega_{*}$, as well as the location of the peak of $\omega_{*}$ in figure 9. Consistent with the magnetic fluctuations presented in figure 5 of [41], we observe that the peak of $\omega_{*}$ has an up-chirping behavior in the first 30-40 ms of the inter-ELM cycle, followed by a flattening. Moreover, we observe that the peak of $\omega_{*}$ starts unaligned with the $q=26/5$ surface, before aligning with it later in the inter-ELM cycle.

### v2

@@ 5 Application to DIII-D Pedestal Data
In this section, we focus on applying the developed techniques to fit pedestal profiles from the DIII-D tokamak [luxon_design_2002], discharge #174864. This discharge, which comes from a series of discharges studying the physics of microtearing modes (MTMs) in DIII-D [nelson_time-dependent_2021], has a particularly wide pedestal as well as sufficient time resolution from the measurements to provide some information about the dynamics of the pedestal between edge localized modes (ELM). In section 5.1 we consider profiles depending on the radial coordinate only, while in sections 5.2 and 5.3 we see how the tools can be used to verify the credibility of features in time-dependent profiles.
@@ 5.1 ELM-averaged Fits
Figure 5: GPR inferences for the electron temperature $T_{e}$ as a function of normalized poloidal flux $\psi$ for DIII-D shot #174864. (Top and middle rows) Data (blue scatter) are shown with corresponding inferred profiles (orange with $\pm 34\%$ confidence interval shaded). The inverse filter cutoff $1/\xi_{*}$ is also shown. Comparing to the width of the pedestal, the right-most profile shows signs of over-regularization. (Bottom row) $N_{eff}$ for the derivative profiles is shown as a function of $\psi$, with $N_{eff}\lesssim 2$ indicating overfitting.
@@
We begin by considering the problem of inferring the profile of the electron temperature $T_{e}$ as a function of the normalized poloidal flux $\psi$ from measurements from the Thomson scattering system [eldon_initial_2012]. Data are aggregated from $N_{t}=44$ time slices of Thomson data taken over 200ms, which cover ELM phases from 5% to 90% over two ELM cycles observed from 2.8s to 3s in the discharge. The data, as well as GPR inferences for three different values of hyperparameters, are shown in figure 5.
@@
The leftmost column displays an inference based on hyperparameters $\sigma^{2},\ell$ set to their MLE values while keeping $\nu=2.5$ fixed. Some ‘wiggles’ appear in the inferred derivative profile, suggesting the possibility of over-fitting. To quantify this, we consider $N_{eff}(\psi)=N_{\mathbf{Y}}[T_{e}^{\prime}(\psi)]$, which represents the effective number of measurements contributing to the inference of the profile derivative $T_{e}^{\prime}$ at a given radial location $\psi$. The bottom row of figure 5 plots this quantity normalized by the number of time slices $N_{t}$. The resulting normalized metric $N_{eff}(\psi)/N_{t}$ provides the effective number of measurements per time slice used to infer the derivative at $\psi$. $N_{eff}/N_{t}\lesssim 2$ indicates that the derivative is primarily inferred using ‘sub-channel’ resolution arising from the radial scatter of the measurements over different time slices. This radial scatter introduces potential covariances in the data that may not be fully captured by the assumed model for the measurement error $\varepsilon_{i}$, suggesting the MLE inference could be over-fitting fluctuations in the data in these regions.
Figure 6: Filled contour plot of the log likelihood $\operatorname{log}p(\mathbf{Y}|\boldsymbol{\theta})$ for various values of $\ell,\sigma^{2}$. The MLE estimate, along with the cases with increased $\ell$, are shown in red crosses. Overplotted on the likelihood contours are contour lines showing the median value of $N_{eff}/N_{t}$ over the strong gradient region, as well as the low-pass filter frequency multiplied by the pedestal width $\xi_{*}w_{ped}$. Dashed contour lines indicate inadmissible hyperparameter values either due to over-fitting or over-regularization.
@@
Figure 7: The top two plots show how varying the length scale $\ell$ hyperparameter changes the inferred values for the peak gradient $\operatorname{max}|dT_{e}/d\psi|$ and full width at half maximum (FWHM) $w_{ped}$ of the temperature pedestal. The bottom plot shows how the median $N_{eff}/N_{t}$ (blue) and the cutoff frequency normalized to pedestal width $\xi_{*}w_{ped}$ (green) change as $\ell$ is varied. The dashed line marks $N_{eff}/N_{t}=3$ and $\xi_{*}w_{ped}=1$.
To explore the hyperparameter space relative to the MLE value, we plot the log likelihood $\log p(\mathbf{Y}|\boldsymbol{\theta})$ in figure 6. Overplotted in blue are contours of $N_{eff}/N_{t}$, which show that as $\ell$ is increased, the risk of over-fitting decreases. Exploring other possible hyperparameter values, the cetner and rightmost columns of figure 5 show inferences with increased $\ell$ relative to the MLE case. These values could plausibly arise from a weighted objective function or hyperprior which biases the hyperparmeter towards larger length scale parameters.
@@
For both of these inferences, $N_{eff}(\psi)/N_{t}\gtrsim 4$ in the steep gradient region, indicating that the inferred derivative values are using information over multiple measurement channels. This reduces concerns about overfitting sub-channel fluctuations in the data associated with the radial scatter. However, there are significant quantitative differences in the inferred profiles for these two cases. The rightmost case appears much smoother, and infers peak gradients that are 8% smaller along with a pedestal width $w_{ped}$, as measured by the full width at half maximum (FWHM) of the gradient profile, that is 13% larger compared to the center case.
@@
In order to systematically quantify the level of regularization applied by GPR, we turn to the low-pass analogy. We consider an effective $\overline{\Delta\psi}=0.2/N_{[0.8,1.0]}$, where $N_{[0.8,1.0]}$ is the total number of measurements of the interval $0.8\leq\psi\leq 1.0$, and $\overline{\sigma_{\varepsilon}^{2}}$ given by the average squared error of the data. We can then compute an effective cutoff frequency $\xi_{*}$ by using (8) with the average signal-to-noise rate $\bar{S}=\sigma^{2}\ell/\overline{\sigma_{\varepsilon}^{2}}\overline{\Delta\psi}$. We plot $\xi_{*}w_{ped}$ in figure 6, showing that the centermost inference satisfies $\xi_{*}w_{ped}\gtrsim 1$, whereas the right-most inference does not.
In addition to plotting the inverse cutoff frequency $1/\xi_{*}$ in figure 5, we also show how the peak gradients, pedestal FWHM, $N_{eff}/N_{t}$, and $\xi_{*}w_{ped}$ systematically change as we vary $\ell$ in figure 7. When $\xi_{*}w_{ped}$ becomes less than unity, the inferred peak gradient decreases and pedestal width increases. This suggests that physical frequencies associated with the pedestal structure are being filtered out by the inference for large values of $\ell$, indicating over-regularization. Meanwhile, when $N_{eff}/N_{t}$ approaches 2, both the inferred peak gradient and the uncertainty in the pedestal width increase. This reflects the inclusion of noise-sensitive higher frequencies for small values of $\ell$, indicating possible over-fitting. Taken together, the $N_{eff}$ metric and the low-pass filter interpretation lend a high degree of confidence in the credibility of the inference in the center column of figure 5 given this point estimate of the hyperparmeters $\ell$. $N_{eff}/N_{t}\gtrsim 4$ in the peak gradient region provides confidence that the pedestal profile is not over-fit, while the low-pass cutoff frequency $\xi_{*}w_{ped}\gtrsim 1$ provides confidence that the pedestal profile is not over-regularized.
Since the center column inference relies on a point estimate of the hyperparameters, it is natural to ask whether it provides an adequate representation of uncertainty compared to a full Bayesian marginalization. Referring to the law of total variance (5), the variation in the inferred quantities as $\ell$ varies in figure 7 gives a simple estimate of how the error bars would change using marginalization. For $0.123\leq\ell\leq 0.217$, these variations remain within the error bars computed from point estimates, indicating that epistemic uncertainty will not dominate over aleatoric uncertainty in this range.
However, it is important to emphasize that the inference $\hat{f}$ represents an approximately low-pass filtered version of $f$. Therefore, it cannot be used to draw conclusions about the presence or absence of features at higher frequencies. The $N_{eff}$ metric and filter cutoff indicate that the available measurements are only sufficient to discriminate pedestal structure within a narrow range of scales. For $\ell>0.217$ the hypotheses are too restrictive on the typical scales for $f$, resulting in over-regularization that smooths away pedestal structure. For $\ell<0.123$ the hypotheses are overly permissive and risk over-fitting. The small epistemic uncertainty thus reflects the limited range of admissible hypotheses on the frequency content of $f$ which are both consistent with the spatial resolution requirements to resolve pedestal structure, and robust against over-fitting. Reminiscent of the ‘coastline paradox’, the peak gradients could in principle become arbitrarily large as higher frequency fluctuations are included, but such possibilities are explicitly filtered out by the inference. With these caveats in mind, and recognizing that unmodeled systematic errors can also increase the size of the error bars, the inference in the center column of figure 5 can be taken as a credible interpretation of the experimental measurements that provides an estimate of the low-pass filtered version of the pedestal profile.
@@ 5.2 Inter-ELM Variation: Setup
To accomplish this, we aggregate measurements over $N_{t}=98$ time slices covering 5 ELM cycles from 2.8s to 3.7s in the discharge. Data from ELMs with irregular frequency during this interval were discarded. We assume that the value of a measurement $Y_{i}$ depends on both the spatial and temporal location of its measurement, $x_{i}=(\psi_{i},t_{i})$. To perform the conditional average, we take $t_{i}$ to be time since last ELM of the measurement. Thus, we are trying to infer a two-dimensional function $p(\psi,t)$ from noisy measurements $Y_{i}=p(\psi_{i},t_{i})+\varepsilon_{i}$. We take a GP prior on $p$ with zero mean, and a covariance which is the direct product of two kernels,
@@
We now use the tools developed in the previous sections to get a point estimate for a set of hyperparameters which provide a credible inference of the inter-ELM evolution. Again, we emphasize that the same caveats discussed before still apply here. Credibility should be understood in terms of giving a point estimate of a low-pass filtered version of the inter-ELM evolution meeting the resolution requirements. Moreover, any error bars need to be understood as representing only the aleatoric uncertainty of this point estimate.
@@
Since we are fitting a pedestal profile, we use $\nu_{\psi}=2.5$. Also, to capture sharp transitions in temporal behavior, we use $\nu_{t}=2.5$. We get an estimate for $\sigma^{2}$ by taking the mean square of the data $Y_{i}^{2}$. This leaves the two length scale parameters $\ell_{\psi},\ell_{t}$ to be determined. In figure 8, we use equation (9) to plot the transfer function $H(\xi,f)$ associated with the product kernel (20) over spatial $\xi$ and temporal $f$ frequencies for a representative choice of $\ell_{\psi}$ and $\ell_{t}$. Notice that the shape of the pass band enclosed by the $H(\xi,f)=1/2$ contour indicates that high temporal frequencies experience stronger attenuation in spatial frequencies, and vice versa.
@@
Next, we average the values of $N_{eff}$ over the spatial range $[0.8,1.0]$ and temporal range $t_{min},t_{max}]$ of the data to compute average values $\bar{N}_{\mathbf{T}}$ and $\bar{N}_{\psi}$. These are used to define the effective spatial and temporal signal-to-noise rates
@@
By keeping the hyperparameters $\sigma^{2},\nu_{\psi},\nu_{t}$ fixed, we can plot $\bar{N}_{\mathbf{T}},\bar{N}_{\psi},\xi_{*}$ and $f_{*}$ as functions of $\ell_{\psi},\ell_{t}$. This is done in figure 9. The ELM period was about $\sim 100$ ms in the considered time range, so a temporal cutoff frequency $f_{*}=30$ cycles per second was chosen, so phenomena with a timescale of $1/f_{*}\approx 33$ ms could be resolved. We remark that without a clear understanding of the physical mechanisms controlling the inter-ELM temporal evolution, it is not possible to conclude if this timescale is sufficient to resolve all of the relevant physical timescales. However, since $f_{*}$ gives an estimate of the temporal smoothing applied by GPR, it is possible to conclude that the inference does not over-regularize on the macroscopic $\sim 100$ ms ELM period timescale. The spatial frequency $\xi_{*}$ was chosen so that the product with the pedestal FWHM $w_{ped}$ would satisfy $\xi_{*}w_{ped}=1$.
@@ 5.3 Inter-ELM Variation: Results
Figure 10: Time-dependent GPR inferences performed using the conditional averaging technique for the electron pressure profile $p_{e}=n_{e}T_{e}$ as a function of $\psi$. (Top and middle rows) Data (scatter) are shown with inferred profile and derivatives $dp_{e}/d\psi$ (curves with $\pm 34\%$ confidence intervals shaded), colored by the time since ELM. The inverse filter cutoff $1/\xi_{*}$ is shown along with low-order rational $q$ surfaces on the derivative profile. (Bottom row) $N_{\psi}$, averaged over $t$, is plotted as a function of $\psi$. $N_{\psi}\lesssim 2$ indicates possible overfitting due to reliance on two or less measurement channels per timeslice for the inference of the derivative.
@@
Recent work has shown how the alignment between the peak of the profile of electron diamagnetic frequency $\omega_{*}$ with low-order rational surfaces in the pedestal can be used to predict features in experimentally observed magnetic fluctuation spectrograms [hatch_microtearing_2021, larakers_global_2021, hassan_identifying_2022, curie_survey_2022, curie_gyrokinetic_2022]. Computing profile gradients is crucial to such an analysis, as $\omega_{*}$ is related to pressure gradients via
@@
Furthermore, following the analysis [nelson_time-dependent_2021], we are able to compare these pressure profiles against the locations of low-order rational $q$ surfaces. Namely, kinetic equilibria are produced every $20$ ms using the CAKE code [xing_cake_2021], then sorted by the time since ELM. The locations of low-order rational $q$ surfaces are computed by averaging over their locations determined from the individual equilibria, and errorbars are computed using the standard error of the mean.
@@
The second interesting feature is the inward motion of the peak of the pedestal pressure gradient as the pedestal profile evolves during the inter-ELM cycle. In [nelson_time-dependent_2021], this behavior was linked to up-chirping of magnetic fluctuations observed in the pedestal region through the predicted dependence of the MTM frequency on the peak value of $\omega_{*}$. We repeat the time-dependent MTM analysis using the profiles inferred from GPR. We plot the evolution of the peak of $T_{e}\frac{\partial(\log p_{e})}{\partial\psi}\propto\omega_{*}$, as well as the location of the peak of $\omega_{*}$ in figure 11. Consistent with the magnetic fluctuations presented in figure 5 of [nelson_time-dependent_2021], we observe that the peak of $\omega_{*}$ has an up-chirping behavior in the first 30-40 ms of the inter-ELM cycle, followed by a flattening. Moreover, we observe that the peak of $\omega_{*}$ starts unaligned with the $q=26/5$ surface, before aligning with it later in the inter-ELM cycle.